OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. 3x - 2y = 9 and -3x + 2y = -9 (3, 0) (5, 6) (1, -3) no solution infinitely many solutions

OpenStudy (anonymous):

Multiplication isn't necessary, just add the two equations together.

OpenStudy (johnweldon1993):

If you combine these 2 equations (elimination method) so add them together... 3x - 2y = 9 -3x + 2y = -9 -------------- 0 Looks like they are the SAME equation...so it looks like there are an infinite amount of solutions